# FLAF: Focal Line and Feature-constrained Active View Planning for Visual Teach and Repeat

### v1

@@ I Introduction
In this paper we present a visual navigation system with the PTU and a novel active view planning method for feature-based VT&R. The contributions of our work are summarized as follows: (1) We combine VSLAM, motion control in repeat, and feature-based active view planning to overcome the low-texture problem. To the best of our knowledge, this is the first demonstration of VT&R system coupled with active view planning. (2) In real-scene experiments of VT&R, we find the active view planner should be redesigned in the repeat phase because of the unidentifiable points in the already existing complete map. We identify that the focal line vector is key to evaluating the angles of a rotatable camera. (3) We propose an online focal line and feature (FLAF)-constrained active view planning method that rewards the camera orientations that observe more feature identifiable map points at their mean-view directions and is suitable for both phases of VT&R.
@@ II-C Active View Planning for VSLAM
The idea to consider both the angles of $\alpha_{1}$ and $\alpha_{2}$ shown in Fig. 3 was first described in the VSLAM system of[9] for feature selection, which serves as the foundation for our system. This strategy is also utilized in the active SLAM method presented in [12], which employs a fixed camera for both navigation and exploration. The model in [12] defines specific ranges of distance and $\alpha_{2}$ to screen map points within the camera’s frustum. The ”FLAF without scoring” method in our comparison experiments can be seen as an implementation of active VT&R using the model from [12], despite the differences in the robot and task. It is noteworthy that Mostegel et al.[25] identified several metrics for feature recognition and validated the effectiveness of using a cosine function to measure the feature recognition probability, which indicates the likelihood of feature identification from different view angles.
@@ III Approach
We build our VT&R system on an active camera-based VSLAM framework shown in Fig. 4, which integrates ORB-SLAM2[9] and our FLAF-constrained active view planning. Our method for active camera-based path following is similar to that described in [17]. During the construction of this navigation system, we observed frequent failures in the repeat phase when feature identifiability was not considered. The core idea of our formulation is to rotate the camera toward optimal directions relative to the map points, thereby improving the performance of the feature detector by taking into consideration the feature identifiability of these map points.
To increase the number of map points in the central area of the field of view, we designed a scoring function based on $\cos({\alpha_{1}})$ to rotate the camera toward regions dense with map points. Each map point within the FoV contributes to the total of this scoring function. Map points with smaller $\alpha_{1}$ values receive higher scores, which encourages the clustering of map points in the central area of the field of view. Consequently, camera poses that capture more map points with smaller $\alpha_{1}$ in their FoV achieve higher scores. This strategy is similar to the UDVP model[16], which uses a sigmoid function with higher computation complexity.
Another goal of our design is to score the map points with good feature identifiability relative to the current camera pose. Feature identifiability refers to the ability of the feature detector to recognize the map point from a specific position and orientation. With the path map built, the VT&R system can determine the existence of a map point. But if the robot rotates the camera toward this map point from an arbitrary position, there is a probability[25] that this map point may not be detected in the current image due to affine changes[10]. In the teach phase, only those map points observed and associated with several different keyframes are retained in the map. We assume the keyframes that observe the same map point define a view distribution within which this map point can be identified. The mean viewing direction is considered the normal of the map point, around which the view distribution is defined. This definition of the map point normal was also used in [9, 12]. Regarding the treatment of the normal of a feature, Mostegel et al.[25] justified using $\cos{\alpha_{2}}$ as the metric of the probability of feature identification to account for the observation of a feature from different viewpoints and view angles. Our scoring function, $\cos({\alpha_{1}})\cdot\cos({\alpha_{2}})$, multiplies these two cosine functions to prioritize map points that score highly on both metrics.
@@ III-A VSLAM with Active Camera
While the mobile robot is moving on a path, images with equally spaced timestamps are sequentially captured by the moving camera and sent to the VSLAM system. For every input image $\mathbf{I}_{k}:\mathbb{R}^{2}\rightarrow{\mathbb{R}}$, ORB features[21] are extracted and sent to the local-map tracking module, in which features are aligned with the local map to estimate current camera pose $\mathbf{X}_{k,w}\in\rm{SE(3)}$ in world coordinate. This process of computing a camera pose according to the local map is called local map tracking. If enough new features are found, this frame is decided as a keyframe and these features are triangulated into the map based on data association between local keyframes. This process of projecting new keyframes and features into the map space (world coordinate) is called local mapping.
@@
For any map point $\mathbf{P}_{j}^{i}$ in $\boldsymbol{M}_{l}$ visible frustum of current camera pose can be matched with an ORB feature at $\mathbf{p}_{j}^{i}\in\mathbb{R}^{2}$ on $\mathbf{I}_{k}$, we reproject it onto the $\mathbf{I}_{k}$ by pinhole camera model $\pi:\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$ and get the reprojected coordinate $\pi(\mathbf{P}_{j}^{i})\in\mathbb{R}^{2}$. Because of the mapping and localization uncertainty, there exists an error $\mathbf{e}_{i,j}\in\mathbb{R}^{2}$ of the j-th map point in i-th keyframe:
@@
But not all points of $\boldsymbol{M}_{l}$ are within visible frustum of the camera and can be matched with an ORB feature in $\mathbf{I}_{k}$. We collect any map point $\mathbf{P}_{j}^{i}$ in $\boldsymbol{M}_{l}$ which can be identified by current view as a set $\boldsymbol{S}=\big{\{}\mathbf{P}_{j}^{i}|\mathbf{P}_{j}^{i}$ can be identified by $\mathbf{I}_{k}\big{\}}$. Then we conduct the local map tracking by minimizing the cost function[9]:
@@
where $\mathbf{X}_{k}^{*}\in\rm{SE}(3)$ represents the pose estimation of $\mathbf{I}_{k}$. The magnitude of map points within the visible frustum of the camera that can be matched with an ORB feature in $\mathbf{I}_{k}$ is represented as:
@@
to indicate that $N_{S}$ is a function of current camera pose and the local map.
According to the results in [12], the pose estimation shown by Equation (3) fails if the magnitude of $S$ is below a threshold. Therefore, [16] proposes a view planner to increase $N_{S}$. To achieve this idea by our sampling-based optimization, the pan-tilt angles of the PTU are sampled as $\mathbf{q}=(pan,tilt)$ and transformed into $\mathbf{T}_{pt}(\mathbf{q})\in\rm{SE}(3)$ to rotate the camera to $\mathbf{X}_{k}^{\prime}$, which is directly evaluated:
$\mathbf{X}_{k}^{\prime}=\mathbf{T}_{pt}(\mathbf{q})\ast\mathbf{X}_{k}$ (5)
III-B Active View Planning for Feature-based VT&R
In the classical feature-based VSLAM[9], on which we build our active camera-based SLAM system, the next frame is typically decided by the trajectory of the robot. To avoid low-texture views, we make passive VSLAM tightly coupled with a feature-based active view planning module. Specifically, we insert (Fig. 4) the active view planning module between the local map tracking and the local mapping because they are separated into different threads in implementation. After local map tracking, we have the local map and current camera pose needed in active view planning. Based on the local map and current camera pose, the next best angles of the PTU can be decided. Thus we achieve using current partial perception results as feedback to control the sensor and image input. This process reflects the concept called “active vision” in the robotics community.
Using the UDVP observation model[16], we control the active camera to see more and closer map points. Given a map point in the local map, the UDVP[16] model awards the PTU angles with a smaller distance to it and a smaller view angle between the line of sight to it and the camera’s focal line. In the teach phase, the active camera always looks at the just-built partial map. In this case, it’s easy to extract ORB features and match them with those in past keyframes because the just-built map points conform to above criteria of UDVP model. Thus $N_{S}=f(\mathbf{X}_{w})$ is increased and tracking failure probability is decreased compared to passive VSLAM. This effect can be represented by the inequality:
@@ III-B Active View Planning for Feature-based VT&R
However, in the repeat phase, a complete map of the taught path is built. Although there exist more local map points that are close to the optical center and have a small angle between sight line $OP$ and focal line $n_{c}$ ($\alpha_{1}$ shown in Fig. 3), some of the points are not identified by the feature matcher because of a big angle between the focal line and mean view angle of the map point. This situation results in the higher failure rate of repeat. This problem can be summarized as “looking at points visible but not identifiable makes the tracking fail”. We draw Fig. 2 to explain the particular VT&R failure.
To solve this problem, we use the observation model proposed in [9] shown in Fig. 3. One distance and two angles are measured for every pan-tilt sample and map point to support our design principles:
@@ III-C Path Learning and Tracking with Active Camera
When the human guide finishes teaching the path, a complete map consisting of plenty of 3D map points and a graph of keyframes are saved by the system. The keyframes with timestamps and poses contain trajectory information of the active camera on this path. In the repeat phase, the map points are used to construct a local map according to the current pose of the camera. Every time a keyframe is stored in memory, the PTU sample is read from the angle encoder and saved with the same timestamp. In summary, the mobile robot is taught with a feature map and a camera trajectory consisting of keyframes with timestamps and corresponding PTU angles. Then the robot trajectory is computed from PTU samples and camera trajectory by inverse operation of Equation (8):
@@
When the mobile robot is needed to navigate the learned path, the feature map taught in the teach phase is loaded at first. Then PnP algorithm[22] is used to find the initial pose of the robot. Once the robot has found its position in the map, the active camera-based SLAM shown in Fig. 2 is booted up to run local map tracking for current pose $\mathbf{X}_{k,w}$. Then we search the Keyframe poses for the closest reference pose $X_{r,w}\in\rm{SE}(3)$ in front of the robot. Finally the pose error between $\mathbf{X}_{k,w}$ and $\mathbf{X}_{r,w}$ is transferred into a PD controller[23] $\boldsymbol{C}_{pd}$ to compute the current velocity $\boldsymbol{\phi}_{k}$:
$\boldsymbol{\phi}_{k}=\boldsymbol{C}_{pd}(\mathbf{X}_{r,w}-\mathbf{X}_{k,w})$ (9)
To speed up the searching of a reference keyframe, we define a window with a width of 10 that is centered at the last reference keyframe. In the repeat phase, the computer keeps outputting the motion calculated by Equation (9) to the wheels.
@@
All the data in Table I are the average results of 10 repeated experiments. ”CR” means the average completion rate in the repeat stage of VT&R.
The AP-RMSE data of Paths 3 and 4 are relative and lack a definite scale because the ground truths are derived using SfM with images captured
by a monocular camera.
@@
”SR” indicates the success rate of completing the entire path in the repeat phase.
@@
Figure 5: Trajectories of passive VT&R and active VT&R using different view planning methods are shown. Offsets on the Y-Axis are manually added to separate the overlapping teaching and repeating trajectories for better visualization. The teaching trajectories for Path 1 and Path 2 were obtained via ”motion capture” as ground truth, while those for Paths 3 and Path 4 were derived using SfM. Our FLAF-based active VT&R system demonstrates the highest completion rate (CR) across all three paths and can reliably navigate all four paths over multiple loops without failure.
@@ IV Experiments and Discussion
In both phases of VT&R, the active camera runs automatically according to current perception. In the teach phase, the robot runs a path following a human guide and automatically learns the path using the active camera-based VSLAM. In the repeat phase, the robot is placed anywhere near the road in a similar orientation to the taught one. Once started, the robot processes the images input by the active camera to localize and control the robot. It needs to be emphasized that the main purpose of our experiments is to demonstrate repeatable and successful VT&R on the difficult paths with our system. As the trajectory error in the repeat phase are all acceptable, we emphasize the results of completion rate, which indicates that only our FLAF-based Active VT&R can finish all three paths. It is worth noting that the UDVP-based active VT&R always fails in the same position and for the same reason in our repeated experiments.
Experiments are performed on three paths to evaluate our active view planning method and VT&R system. The first two paths are in the effective range of the motion capture device in the Shenzhen Key Laboratory of Robotics and Computer Vision. Path three leads the robot from an indoor location to a reading space out of the laboratory and finally back to the start. Experiments on each of the three paths are performed 10 times for consistent results. All the data in Table I are the average results of 10 repeated experiments. The trajectories shown in Fig. 5 and the point graph shown in Fig. 7 are a representative selection, considering that the traditional methods always fail in the same situation. In Fig. 6, we demonstrate the process of experiments on the most difficult path connecting different rooms.
We compare our FLAF-based active view planning with three other methods on the task of VT&R. ”Passive VT&R” is achieved with our VT&R system without the PTU and active view planning. The ”UDVP-based active VT&R” is a reproduction of the observation model proposed in [16] with our VT&R system. The only difference between FLAF and UDVP is the observation model. The equation(7) is supported by the comparison of ”FLAF-based active” with ”FLAF by counting,” which is the first version of FLAF that has been eliminated by us.
@@ IV-A Implementation and Experimental Setup
As shown in Figure 1(d), we fix an Intel Realsense D435 camera on an I-Quotient-Robotics PTU to make up our active camera. This active camera is mounted on a Clearpath-Jackal robot. Our active VT&R system runs in real-time on a notebook with an Intel i7 (2.3GHz) CPU and responds to the images exactly at the frame rate of 20Hz.
Ground truths of path one and two are obtained from motion capture in the teach phase. The ground truth of Path 3 is built by SfM, as there is no motion capture outside the laboratory. The visual repeating trajectories of all methods are compared with the ground truths and painted in Fig. 5. In Table I, the CR data shown indicates the VT&R completion rate of this view planning method, and the time data indicates the mean time used for the view planning methods implemented by sampling-based optimization. The efficacy of VSLAM is decreased by the relatively low speed of view planning compared to the SLAM speed of 20-30Hz. The translational RMSE of trajectories is computed using evo[24] by comparing the repeat trajectory with the taught one based on timestamps, resulting in greater error than physical truth. The trajectory consistencies between repeat and teach shown in Fig. 4 are similar and all acceptable.
On path one and two, we demonstrate the efficacy of our VT&R system in both an active and a passive way. On path three (Fig. 6), we show a challenging case with low-texture regions where passive VT&R fails and active VT&R succeeds. Additionally, our FLAF model is verified on all three paths to outperform the existing UDVP in repeating a complete path. FLAF by counting refers to counting the map points in a range defined by FLAF instead of grading the points by the product of the cosine functions shown in Equation (7).

### v2

@@ I Introduction
In this paper we present a visual navigation system with the PTU and a novel active view planning method for feature-based VT&R. The contributions of our work are summarized as follows: (1) We combine VSLAM, motion control in repeat, and feature-based active view planning to overcome the low-texture problem. To the best of our knowledge, this is the first demonstration of VT&R system coupled with active view planning. (2) We propose an online focal line and feature (FLAF)-constrained active view planning method that rewards the camera orientations that observe more feature identifiable map points at their mean-view directions and is suitable for both phases of VT&R.
@@ II-C Active View Planning for VSLAM
The idea to consider both the angles of $\alpha_{1}$ and $\alpha_{2}$ shown in Fig. 3 was first described in the VSLAM system of[9] for feature selection, which serves as the foundation for our system. This strategy is also utilized in the active SLAM method presented in [12], which employs a fixed camera for both navigation and exploration. The model in [12] defines specific ranges of distance and $\alpha_{2}$ to screen map points within the camera’s frustum. The “FLAF without scoring” method in our comparison experiments can be seen as an implementation of active VT&R using the model from [12], despite the differences in the robot and task. It is noteworthy that Mostegel et al.[25] identified several metrics for feature recognition and validated the effectiveness of using a cosine function to measure the feature recognition probability, which indicates the likelihood of feature identification from different view angles.
@@ III Approach
We build our VT&R system on the active camera-based VSLAM framework shown in Fig. 4, which integrates ORB-SLAM2[9] and our FLAF-constrained active view planning. Our method for active camera-based path following is similar to that described in [17], which resolves the robot poses from camera localization and PTU angles as the input of VT&R.
To increase the number of map points in the central area of the FoV, we designed a scoring function (Eq. (3)) based on $\cos({\alpha_{1}})$ (see Fig. 3) to rotate the camera toward regions dense with map points. As the UDVP model[16] uses a sigmoid function of $\alpha_{1}$, we apply a cosine function to reduce the computational cost. By adding up the scores of all map points within the FoV, each one of them contributes to the total score of this PTU sample. Map points with smaller $\alpha_{1}$ values receive higher scores, which encourages the clustering of map points in the central area of the FoV. Consequently, samples of the active camera that capture more map points with smaller $\alpha_{1}$ in their FoV achieve higher scores.
Another goal of our design is to score the map points with good feature identifiability relative to the samples of PTU angles. Feature identifiability refers to the ability of the feature detector to recognize the map point from a specific position and orientation. With the path map built, the VT&R system can determine the existence of a map point at a 3D coordinate. But if the robot rotates the camera toward this map point from an arbitrary position, there is a probability[25] that this map point may not be detected in the current image due to affine changes[10]. We assume the keyframes that observe the same map point define a view distribution within which this map point can be identified. The mean viewing direction is considered the normal of the map point, around which the view distribution is defined. This definition of the map point normal was also used in [9, 12]. Mostegel et al.[25] justified using $\cos{\alpha_{2}}$ as the metric of the probability of feature identification to account for the observation of a feature from different viewpoints and view angles. Our scoring function, $\cos({\alpha_{1}})\cdot\cos({\alpha_{2}})$, multiplies two cosine functions to prioritize map points that score highly on both metrics.
@@ III-A VSLAM with Active Camera
In the classical feature-based VSLAM[9], on which we build our active camera-based SLAM system, the next frame is passively decided by the trajectory of the robot. To avoid low-texture views, we make passive VSLAM tightly coupled with a feature-based active view planning module. Specifically, we insert (Fig. 4) the active view planning module between the local map tracking and the local mapping as they are separated into different threads in implementation.
While the mobile robot is moving along a path, images with equally spaced timestamps are sequentially captured by the moving camera and sent to the VSLAM system. For every input image $\mathbf{I}_{k}:\mathbb{R}^{2}\rightarrow{\mathbb{R}}$, ORB features[21] are extracted and sent to the local-map tracking module, in which features are aligned with the local map to estimate current camera pose $\mathbf{X}_{k,w}\in\rm{SE(3)}$ in world coordinate. This process of computing a camera pose according to the local map is called local map tracking. If enough new features are found, this frame is decided as a keyframe and these features are triangulated into the map based on data association between local keyframes. This process of projecting new keyframes and features into the map space (world coordinate) is called local mapping.
@@
For any $\mathbf{P}_{j}^{i}$ in $\boldsymbol{M}_{l}$ that is within the visible frustum of the current camera pose, if it can be matched with an ORB feature at $\mathbf{p}_{j}^{i}\in\mathbb{R}^{2}$ on $\mathbf{I}_{k}$, it is reprojected onto the $\mathbf{I}_{k}$ by pinhole camera model $\pi:\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$ and get the reprojected coordinate $\pi(\mathbf{P}_{j}^{i})\in\mathbb{R}^{2}$. The reprojection error as in Eq. (2) is applied for optimizations in VSLAM:
@@
We use a set $\boldsymbol{S}=\big{\{}\mathbf{P}_{j}^{i}|\mathbf{P}_{j}^{i}$ can be identified by $\mathbf{I}_{k}\big{\}}$t o represent the map points in $\boldsymbol{M}_{l}$ that can be identified by the current view. Then local map tracking is conducted by minimizing the cost function[9]:
@@
where $\mathbf{X}_{k}^{*}\in\rm{SE}(3)$ represents the pose estimation of $\mathbf{I}_{k}$.
After local map tracking, we have the local map and current camera pose needed in active view planning. Based on the local map and current camera pose, the next best angles of the PTU can be decided. Thus we achieve using current partial perception results as feedback to control the sensor and image input. According to the results in [12], the pose estimation shown by Equation (3) fails if the magnitude of $S$ is below a threshold. Therefore, [16] proposes a view planner to increase $N_{S}$. To achieve this by our sampling-based optimization, the pan-tilt angles are sampled as $\mathbf{q}=(pan,tilt)$ and transformed into $\mathbf{T}_{pt}(\mathbf{q})\in\rm{SE}(3)$ to obtain the corresponding sample of camera pose $\mathbf{X}_{k}^{\prime}$, which is directly evaluated:
$\mathbf{X}_{k}^{\prime}=\mathbf{T}_{pt}(\mathbf{q})\ast\mathbf{X}_{k}$ (4)
while the robot is moving along a path, the PTU samples are evaluated by view planners at a set frequency, and the best one is sent to the PTU to rotate the active camera. During the simultaneous movements of the PTU and the robot, the images input to the VSLAM are decided by the PTU angles and the robot trajectory, while the PTU angles are decided by the robot trajectory and the local map of the VSLAM system.
III-B Active View Planning for Feature-based VT&R
We use Eq. (5) to represent the magnitude of map points within the visible frustum of the camera that can be matched with an ORB feature in $\mathbf{I}_{k}$:
@@ III-B Active View Planning for Feature-based VT&R
Eq. (5) indicates that $N_{S}$ is a function of the current camera pose and the local map. Using the UDVP observation model[16], we control the active camera to see more and closer map points. Given a PTU sample, the UDVP[16] model scores map points with a smaller distance to the camera center and a smaller view angle between the line of sight to them and the camera’s focal line. Thus $N_{S}=f(\mathbf{X}_{w})$ is increased and tracking failure probability is decreased compared to passive VSLAM. This effect can be represented by the inequality:
@@
However, in the repeat phase, a complete map of the taught path is built. Although there exist more local map points that are close to the optical center and have a small angle between $OP$ and $n_{c}$ ($\alpha_{1}$ shown in Fig. 3), some of them can not be identified by the feature detector due to a big angle between the focal line and mean view angle of the map point. This phenomenon can be summarized as “looking at points visible but not identifiable”, which results in a higher failure rate of repeat. To address this, we apply the observation model shown in Fig. 3. One distance and two angles are measured for every PTU sample and map point to support our view planner:
@@ III-C Path Learning and Tracking with Active Camera
After Teaching, a complete map consisting of plenty of 3D map points and a graph of keyframes are saved with corresponding PTU angles read from the angle encoder. The keyframes with timestamps, poses, and PTU samples contain trajectory information of the active camera and the robot. Then the robot trajectory is computed from PTU samples and camera trajectory by inverse operation of Equation (8):
@@
Once the repeat begins, the previously taught feature map is loaded, and the PnP algorithm[22] is executed to determine the initial camera pose. The active camera-based VSLAM, as shown in Fig. 2, is then performed to compute the current pose $\mathbf{X}_{k,w}$. Following this, we search for the closest keyframe pose $X_{r,w}\in\rm{SE}(3)$ ahead of the current localization. Finally, the pose error between $\mathbf{X}_{k,w}$ and $\mathbf{X}_{r,w}$ is processed by a PD controller[23], denoted as $\boldsymbol{C}_{pd}$, to calculate the current velocity $\boldsymbol{\phi}_{k}$:
$\boldsymbol{\phi}{k}=\boldsymbol{C}{pd}(\mathbf{X}{r,w}-\mathbf{X}{k,w})$ (9)
To expedite the reference keyframe search, we define a search window centered on the last reference keyframe, with a fixed width of 10.
@@
[…]
and 4 are relative and lack a definite scale because the ground truths are derived using SfM with images captured by a monocular camera.
“CR” means the average completion rate in the repeat stage of VT&R. “Time” means the average time used by the sampling-based view planner.
@@
“SR” indicates the success rate of completing the entire path in the repeat phase.
@@
Figure 5: Trajectories of passive VT&R and active VT&R using different view planning methods are shown. Offsets on the Y-Axis are manually added to separate the overlapping teaching and repeating trajectories for better visualization. The teaching trajectories for Path 1 and Path 2 were obtained via “motion capture” as ground truth, while those for Paths 3 and Path 4 were derived using SfM. Our FLAF-based active VT&R system demonstrates the highest completion rate (CR) across all three paths and can reliably navigate all four paths over multiple loops without failure.
@@ IV Experiments and Discussion
In both phases of VT&R, the active camera operates automatically according to current perception. During the teaching phase, the robot follows a human-guided path and learns the path automatically using the active camera-based VSLAM. Before the repeating, the robot is placed anywhere near the taught path with a similar orientation to the taught one. Once the repeating begins, the robot autonomously runs along the taught path, adjusting the camera orientation as needed.
The main purpose of our experiments is to demonstrate the repeatable and successful VT&R on the challenging paths with our system. As the trajectory errors of repeating are all acceptable, we emphasize the completion and success rate (CR and SR) as the key results, indicating that only our FLAF-based Active VT&R can complete all four paths at a high SR. Notably, the UDVP-based active VT&R consistently fails at the same location across multiple repetitions.
We compared our FLAF-based active view planning against three other methods: Passive VT&R, UDVP-based active VT&R, and FLAF without scoring-based active VT&R. Passive VT&R is achieved on our VT&R system without incorporating active view planning. The UDVP-based active VT&R reproduced the observation model proposed in [16] with our VT&R framework. The comparison with “FLAF without scoring” serves as an ablation study, illustrating the impact of our scoring mechanism as described by equation (7). FLAF without scoring-based active VT&R can also be seen as an active VT&R with the model depicted in [12].
@@ IV-A Implementation and Experimental Setup
As shown in Figure 1(d), we fix an Intel Realsense D435 camera on an I-Quotient-Robotics PTU to make up our active camera, which is mounted on a Clearpath-Jackal robot. Our active VT&R system runs in real-time on a notebook with an Intel i7 (2.3GHz) CPU and responds to the images exactly at the frame rate of 20Hz.
Experiments are performed on four paths to evaluate our active view planning method and VT&R system. The first two paths are in the effective range of the motion capture device in the Shenzhen Key Laboratory of Robotics and Computer Vision. Paths 3 and 4 respectively lead the robot from inside the laboratory to a space out of the laboratory and finally back to the start. All the data in Table I are the average results of 10 repeated experiments. The trajectories shown in Fig. 5 and the plots shown in Fig. 7 are a representative selection, considering that the previous methods consistently fail in the same situation. In Fig. 6, we demonstrate the process during which our active VT&R system overcomes the low-texture problem on path 3 that connects different rooms.
Ground truths of paths one and two are obtained from motion capture in the teaching phase. The ground truth of Path 3 is built by SfM, as there is no motion capture outside the laboratory. The visual repeating trajectories of all methods are compared with the ground truths and painted in Fig. 5. In Table I, the CR data shown indicates the VT&R completion rate of this view planning method, and the time data indicates the mean time used for the view planning methods implemented by sampling-based optimization. The efficacy of VSLAM is decreased by the relatively low speed of view planning compared to the SLAM speed of 20-30Hz. The translational RMSE of trajectories is computed using evo[24] by comparing the repeat trajectory with the taught one based on timestamps, resulting in greater error than physical truth. The trajectory consistencies between repeat and teach shown in Fig. 4 are all acceptable for T&R.
On paths 1 and 2, we demonstrate the efficacy of our VT&R system in both an active and a passive way. On paths 3 (Fig. 6) and 4, we show challenging cases with low-texture regions where passive VT&R fails and active VT&R succeeds. Additionally, our FLAF model is verified on all four paths to outperform the existing UDVP in repeating a complete path. FLAF without scoring refers to counting the map points in a range defined by FLAF instead of grading the points by the product of the cosine functions shown in Equation (7).